# Project Alexandria: Towards Freeing Scientific Knowledge from Copyright Burdens via LLMs

### v1

@@ Abstract
Paywalls, licenses and copyright rules often restrict the broad dissemination and reuse of scientific knowledge. We take the position that it is both legally and technically feasible to extract the scientific knowledge in scholarly texts. Current methods, like text embeddings, fail to reliably preserve factual content, and simple paraphrasing may not be legally sound. We urge the community to adopt a new idea: convert scholarly documents into Knowledge Units using LLMs. These units use structured data capturing entities, attributes and relationships without stylistic content. We provide evidence that Knowledge Units (1) form a legally defensible framework for sharing knowledge from copyrighted research texts, based on legal analyses of German copyright law and U.S. Fair Use doctrine, and (2) preserve most ($\sim$95%) factual knowledge from original text, measured by MCQ performance on facts from the original copyrighted text across four research domains. Freeing scientific knowledge from copyright promises transformative benefits for scientific research and education by allowing language models to reuse important facts from copyrighted text. To support this, we share open-source tools for converting research documents into Knowledge Units. Overall, our work posits the feasibility of democratizing access to scientific knowledge while respecting copyright.
@@
Ameya Prabhu2,4⋄   Tawsif Ahmed1,2⋄   Andreas Hochlehnert2,4⋄   Huu Nguyen1,2,5⋄   Nick Akinci Heidrich6⋄
@@ 10 Technical University of Munich, Germany
Knowledge-Units Database Project Page Codebase
@@ 2 Knowledge Units
The evolution of the Earth-Moon system is described by the dark matter field fluid model proposed in the Meeting of Division of Particle and Field 2004, American Physical Society. The current behavior of the Earth-Moon system agrees with this model very well, and the general pattern of the evolution of the Moon-Earth system described by this model agrees with geological and fossil evidence. The closest distance of the Moon to Earth was about 259000 km at 4.5 billion years ago, which is far beyond the Roche’s limit. The result suggests that the tidal friction may not be the primary cause for the evolution of the Earth-Moon system. The average dark matter field fluid constant derived from Earth-Moon system data is $4.39\times 10^{-22}\,\text{s}^{-1}\text{m}^{-1}$. This model predicts that Mars’s rotation is also slowing with the angular acceleration rate about $4.38\times 10^{-22}\,\text{rad}\,\text{s}^{-2}$. Context: "The provided text focuses on the style and genre of a scientific paper…"
@@
Comparison to Knowledge Graphs Knowledge Units share similarities with knowledge graphs (Hogan et al., 2021); however, they are generated locally, on a per-paragraph basis, rather than globally. Unlike conventional knowledge graph tools such as REBEL (Huguet Cabot & Navigli, 2021), which decompose text into concise triples (subject, relation, object), Knowledge Units maintain the contextual richness and nuances of the original paragraph, ensuring that relationships and entities remain closely aligned with the source text. However, KUs lack more logic-based knowledge representation such as OWL-based knowledge graphs, and entities in KUs are not equipped with worldwide unique identifiers (URI/IRIs). While these are significant disadvantages for general-purpose knowledge representation, it is not required for our prime use case of knowledge and context provision for downstream AI applications.
@@ 3 Legal Considerations for Fair Knowledge Extraction
We start our legal considerations under the broader legal framework of the Universal Declaration of Human Rights (UDHR), Article 27(a) which states “[e]veryone has the right freely to participate in the cultural life of the community, to enjoy the arts and to share in scientific advancement and its benefits.” However, it also specifies the need to support and respect the rights of authors in their “moral and material interests resulting from any scientific, literary or artistic production of which (s)he is the author” as also stated under the UDHR 27(b). These rights of authors must co-exist with everyone’s rights to scientific knowledge, and indeed freeing such knowledge from copyright burdens will allow scholars to participate in the marketplace of ideas - to continue new research. From this starting point, we analyzed in detail the legal implications of extracting knowledge from copyrighted text under two legal frameworks: German and U.S. copyright law, presented below.
@@ 4.3 Results: Assessing Content Overlap
Q1: Is there a significant overlap? The top portion of Table 4 reports 5-gram, 7-gram, and 11-gram Jaccard similarities for the entire dataset, as well as for the top 5% of the most similar original text–KU pairs. The Gemini-1.5 Flash model consistently exhibits very low overlap (<<3% in the most conservative 5-gram scenario), even in the top 5% subset. The Qwen-2.5 model shows slightly higher scores but remains below 7% in the same scenario. Overall, these findings indicate negligible direct textual reuse.
@@ 5.2 Addressing Common Criticisms
Rebuttal: We agree. We try to mitigate this by careful design, using structured extraction frameworks converting text into non-expressive factual units.
@@
Rebuttal: We agree, and believe a critical next step is designing hybrid systems integrating cross-referencing algorithms, confidence scoring, and targeted human oversight for critical assertions to balance scalability with rigor. Moreover, we believe that upcoming more capable models, like recently developed reasoning models, will provide further way to automate and scale verification of extracted knowledge content, requiring human assistance only in few manageable cases.
@@
Rebuttal: Our proposal presents an idea with a prototype to demonstrate feasibility – this is a position paper rather than a proposed system. We stress the need for further iteration before deployment and advocate for exploratory dialogue over immediate adoption to prevent uncritical implementation.
@@ Acknowledgements
The authors would like to thank (in alphabetical order): Sebastian Dziadzio, Kristof Meding, Tea Mustać, Shantanu Prabhat for insightful feedback and suggestions. Special thanks to Andrej Radonjic for help in scaling up data collection. GR and SA acknowledge financial support by the German Research Foundation (DFG) for the NFDI4DataScience Initiative (project number 460234259). AP and MB acknowledge financial support by the Federal Ministry of Education and Research (BMBF), FKZ: 011524085B and Open Philanthropy Foundation funded by the Good Ventures Foundation. AH acknowledges financial support by the Federal Ministry of Education and Research (BMBF), FKZ: 01IS24079A and the Carl Zeiss Foundation through the project ”Certification and Foundations of Safe ML Systems” as well as the support from the International Max Planck Research School for Intelligent Systems (IMPRS-IS). JJ acknowledges funding by the Federal Ministry of Education and Research of Germany (BMBF) under grant no. 01IS22094B (WestAI - AI Service Center West), under grant no. 01IS24085C (OPENHAFM) and under the grant DE002571 (MINERVA), as well as co-funding by EU from EuroHPC Joint Undertaking programm under grant no. 101182737 (MINERVA) and from Digital Europe Programme under grant no. 101195233 (openEuroLLM)
@@ B.1 Abstract-Level Analysis
The evolution of the Earth-Moon system is described by the dark matter field fluid model proposed in the Meeting of Division of Particle and Field 2004, American Physical Society. The current behavior of the Earth-Moon system agrees with this model very well and the general pattern of the evolution of the Moon-Earth system described by this model agrees with geological and fossil evidence. The closest distance of the Moon to Earth was about 259000 km at 4.5 billion years ago, which is far beyond the Roche’s limit. The result suggests that the tidal friction may not be the primary cause for the evolution of the Earth-Moon system. The average dark matter field fluid constant derived from Earth-Moon system data is $4.39x10^{(}-22)s^{(}-1)m^{(}-1)$. This model predicts that the Mars’s rotation is also slowing with the angular acceleration rate about $4.38x10^{(}-22)rads^{(}-2)$. What is the primary implication of the dark matter field fluid model’s agreement with the current Earth-Moon system behavior and geological evidence?

### v2

@@ Abstract
Paywalls, licenses and copyright rules often restrict the broad dissemination and reuse of scientific knowledge. We take the position that it is both legally and technically feasible to extract the scientific knowledge in scholarly texts. Current methods, like text embeddings, fail to reliably preserve factual content, and simple paraphrasing may not be legally sound. We propose a new idea for the community to adopt: convert scholarly documents into knowledge preserving, but style agnostic representations we term Knowledge Units using LLMs. These units use structured data capturing entities, attributes and relationships without stylistic content. We provide evidence that Knowledge Units (1) form a legally defensible framework for sharing knowledge from copyrighted research texts, based on legal analyses of German copyright law and U.S. Fair Use doctrine, and (2) preserve most ($\sim$95%) factual knowledge from original text, measured by MCQ performance on facts from the original copyrighted text across four research domains. Freeing scientific knowledge from copyright promises transformative benefits for scientific research and education by allowing language models to reuse important facts from copyrighted text. To support this, we share open-source tools for converting research documents into Knowledge Units. Overall, our work posits the feasibility of democratizing access to scientific knowledge while respecting copyright.
@@
Ameya Prabhu2,4⋄  Tawsif Ahmed1,2⋄  Andreas Hochlehnert2,4⋄  Huu Nguyen1,2,5⋄  Nick Akinci6⋄
@@ 10 Technical University of Munich, Germany
[…]
@@ 2 Knowledge Units
The evolution of the Earth-Moon system is described by the dark matter field fluid model proposed in the Meeting of Division of Particle and Field 2004, American Physical Society. The current behavior of the Earth-Moon system agrees with this model very well, and the general pattern of the evolution of the Moon-Earth system described by this model agrees with geological and fossil evidence. The closest distance of the Moon to Earth was about 259000 km at 4.5 billion years ago, which is far beyond the Roche’s limit. The result suggests that the tidal friction may not be the primary cause for the evolution of the Earth-Moon system. The average dark matter field fluid constant derived from Earth-Moon system data is $4.39\times 10^{-22}\,\text{s}^{-1}\text{m}^{-1}$. This model predicts that Mars’s rotation is also slowing with the angular acceleration rate about $-4.38\times 10^{-22}\,\text{rad}\,\text{s}^{-2}$. Context: "The text appears in a scientific discussion on how the Earth-Moon system’s evolution can be explained by a dark matter field fluid model…"
@@
Comparison to Knowledge Graphs Knowledge Units are very similar to knowledge graphs (Hogan et al., 2021); however, they are generated locally, on a per-paragraph basis, rather than globally. One primary difference is that Knowledge Graphs require to have either locally or even globally unique identifiers (URI/IRIs). This is not the case for the KU graphlets we extract - for two KUs extracted from neighboring paragraphs of a source document, the entities and relationships might differ, even when referring to the same concepts. Hence, KUs lack more logic-based knowledge representation such as OWL-based knowledge graphs. Secondly, unlike conventional knowledge graph tools such as REBEL (Huguet Cabot & Navigli, 2021), which decompose text into concise triples (subject, relation, object), Knowledge Units maintain the contextual richness and nuances of the original paragraph, ensuring that relationships and entities remain closely aligned with the source text. While these are significant disadvantages for general-purpose knowledge representation, it is not required for our prime use case of knowledge and context provision for downstream AI applications.
@@ 3 Legal Considerations for Fair Knowledge Extraction
We start our legal considerations under the broader human rights framework of the Universal Declaration of Human Rights (UDHR), Article 27(a) which states “[e]veryone has the right freely to participate in the cultural life of the community, to enjoy the arts and to share in scientific advancement and its benefits.” However, it also specifies the need to support and respect the rights of authors in their “moral and material interests resulting from any scientific, literary or artistic production of which (s)he is the author” as stated under the UDHR 27(b). These rights of authors must co-exist with everyone’s rights to scientific knowledge, and indeed freeing such knowledge from copyright burdens will allow scholars to participate in the marketplace of ideas - to continue new research. From this starting point, we analyzed in detail the legal implications of extracting knowledge from copyrighted text under two legal frameworks: German and U.S. copyright law, presented below.
@@ 4.3 Results: Assessing Content Overlap
Q1: Is there a significant overlap? The top portion of Table 4 reports 5-gram, 7-gram, and 11-gram Jaccard similarities for the entire dataset, as well as for the top 5% of the most similar original text–KU pairs. The Gemini-1.5 Flash model consistently exhibits very low overlap ($<$3% in the most conservative 5-gram scenario), even in the top 5% subset. The Qwen-2.5 model shows slightly higher scores but remains below 7% in the same scenario. Overall, these findings indicate negligible direct textual reuse.
@@ 5.2 Addressing Common Criticisms
Rebuttal: We agree. We try to mitigate this by careful design, using structured extraction frameworks converting text into non-expressive factual units. We additionally highlight the process must involve only temporary access to original copyrighted texts, strictly for the purpose of analysis under TDM exceptions, followed by their deletion. No original copyrighted material should be retained.
@@
Rebuttal: We agree, and believe a critical next step is designing hybrid systems integrating cross-referencing algorithms, confidence scoring, and targeted human oversight for critical assertions to balance scalability with rigor. The KUs would be akin to editing a machine-readable Wikipedia (human-readable with good UI), one can simply edit the incorrect KU and replace the wrong information. Moreover, we believe that upcoming more capable models, like recently developed reasoning models, will provide further way to automate and scale verification of extracted knowledge content, requiring human assistance only in few manageable cases.
@@
Rebuttal: Our proposal presents an idea with a prototype to demonstrate feasibility – this is a position paper rather than a proposed system. We stress the need for further iteration before deployment and advocate for exploration over immediate adoption to prevent uncritical implementation. Other measures, designed to reduce hallucination propagation can similarly update KU contents in this case, mitigating any harm caused similar to those adopted for leaks of copyrighted data.
6. Replication of failures of Semantic Web. We think semantic web initiatives such as schema.org or life-science KGs have excelled in adoption and impact. However, the standardization done in fully fledged KGs is not (yet) feasible (e.g., wrt. unique identification and schema harmonization) and hence omitted in the design of KUs to avoid past failures, as it is not critical for our purpose. We want to enable KUs to primarily allow sharing of knowledge on platforms like HuggingFace contribute strongly to open science.
@@ Acknowledgements
The authors would like to thank (in alphabetical order): Sebastian Dziadzio, Kristof Meding, Tea Mustać, Shantanu Prabhat for insightful feedback and suggestions. We thank the Grass Foundation and Andrej Radonjic, with special thanks to Andrii Prolorenzo of Wynd Labs for help in scaling up data collection. GR and SA acknowledge financial support by the German Research Foundation (DFG) for the NFDI4DataScience Initiative (project number 460234259). AP and MB acknowledge financial support by the Federal Ministry of Education and Research (BMBF), FKZ: 011524085B and Open Philanthropy Foundation funded by the Good Ventures Foundation. AH acknowledges financial support by the Federal Ministry of Education and Research (BMBF), FKZ: 01IS24079A and the Carl Zeiss Foundation through the project ”Certification and Foundations of Safe ML Systems” as well as the support from the International Max Planck Research School for Intelligent Systems (IMPRS-IS). JJ acknowledges funding by the Federal Ministry of Education and Research of Germany (BMBF) under grant no. 01IS22094B (WestAI - AI Service Center West), under grant no. 01IS24085C (OPENHAFM) and under the grant 16HPC117K (MINERVA), as well as co-funding by EU from EuroHPC Joint Undertaking programm under grant no. 101182737 (MINERVA) and from Digital Europe Programme under grant no. 101195233 (openEuroLLM)
@@ B.1 Abstract-Level Analysis
The evolution of the Earth-Moon system is described by the dark matter field fluid model proposed in the Meeting of Division of Particle and Field 2004, American Physical Society. The current behavior of the Earth-Moon system agrees with this model very well and the general pattern of the evolution of the Moon-Earth system described by this model agrees with geological and fossil evidence. The closest distance of the Moon to Earth was about 259000 km at 4.5 billion years ago, which is far beyond the Roche’s limit. The result suggests that the tidal friction may not be the primary cause for the evolution of the Earth-Moon system. The average dark matter field fluid constant derived from Earth-Moon system data is $4.39x10^{(}-22)s^{(}-1)m^{(}-1)$. This model predicts that the Mars’s rotation is also slowing with the angular acceleration rate about $-4.38x10^{(}-22)rads^{(}-2)$. What is the primary implication of the dark matter field fluid model’s agreement with the current Earth-Moon system behavior and geological evidence?